# Segregation at prior austenite grain boundaries: the competition between boron and hydrogen

### v1

@@ Abstract
The interaction between boron and hydrogen at grain boundaries has been investigated experimentally and numerically in boron-doped and boron-free martensitic steels using thermal desorption spectrometry (TDS) and ab initio calculations. The calculations show that boron, which mostly segregates into prior austenite grain boundaries (PAGBs), repels hydrogen. This behavior has also been observed using TDS measurements, with the disappearance of one peak when boron is incorporated into the microstructure. Additionally, the microstructure of both boron-doped and boron-free steels has been studied through electron backscattered diffraction, electron channeling contrast imaging, synchrotron X-ray measurements, and atom probe tomography. While both steels have a similar grain size, grain boundary distribution, and dislocation densities, pronounced boron segregation into PAGBs is observed for boron-doped steels. Therefore, the equilibrium hydrogen concentration in different trapping sites has been evaluated using the Langmuir-McLean approximation. This thermodynamic model shows that the distribution of hydrogen is identical for all traps when the total hydrogen concentration is low for boron-free steel. However, when it increases, traps of the lowest segregation energies (mostly PAGBs) are firstly saturated, which promotes failure initiation at this defect type. This finding partially explains why PAGBs are the weakest microstructure feature when martensitic steels are exposed to hydrogen-containing environments.
@@ 1 Introduction
Hydrogen is a promising carrier molecule for a low-carbon economy and many corresponding industrial projects are under development to use it as an energy carrier or to develop sustainable processes like the direct reduction of iron ore [1, 2, 3, 4]. It is also known that hydrogen easily diffuses in metals and alloys inducing premature failure of engineering structures, a phenomenon referred to as hydrogen embrittlement (HE). HE is well-known and has been described by numerous models in the literature and thoroughly reviewed in refs. [5, 6]. However, these models provide only a phenomenological description of HE (e.g.: hydrogen enhancing the formation of vacancies [7, 8, 9], localizing plasticity [10, 11, 12], reducing the cohesive interface energy of the system [13, 14] and so on) because the direct observation of hydrogen at the atomic scale is quite challenging given the difficulty to locate hydrogen in alloys experimentally [15, 16, 17].
@@
Steel is one of the candidate materials envisaged to contain and transport hydrogen due to its high mechanical strength, low price, flexibility to form different microstructures from its chemical composition, and the variety of processes that can be applied to the alloys [18]. However, it is also subject to HE [19]. Different processes have been developed to reduce HE of steels [20, 21], including reducing the grain size [22], which is efficient up to a limit, and also modifies other properties of the alloy, potentially to a point of being unsuitable for the envisaged application. Steel composition features have also been modified, for instance, to form precipitates and reduce the negative impact of hydrogen [23, 24]. However, once the traps inherent to the precipitates, including interfaces with the matrix or sites inside the precipitates, are saturated with hydrogen, they become inefficient for further HE mitigation specifically in typical engineering loading scenarios, which are often characterized by permanent hydrogen exposure. More recently, the implementation of chemically heterogeneous Mn profiles in the austenite phase of medium Mn-steels has also been proposed, improving the ductility of the alloys in hydrogen [25]. However, such treatment is only possible for a relatively narrow range of Mn-steel compositions. Finally, another recently proposed solution to mitigate HE in martensitic steels is the segregation of interstitial solutes to structural defects, being the weakest microstructure features in hydrogen environments [26]. In the case of martensitic steels, these weakest features are prior austenite grain boundaries (PAGBs) [27, 26]. By favoring the segregation of carbon into these defects using special heat treatments, the negative impact of hydrogen is attenuated [26]. Thus, this work is focused on the effect of boron localized at this structural defect and its impact on the segregation of hydrogen.
@@
Boron is an alloying element often incorporated in steel because it improves the hardenability and delays the formation of ferrite [28]. It enhances the properties of the alloy by segregating into PAGBs of the iron alloy. However, the segregation mechanism of boron remained elusive until recently, when it was described using equilibrium segregation and considering the precipitation of borides or boro-carbides at GBs [29, 30]. It has also been observed recently that boron can improve the resistance to HE of steels [31], which is also observed in other alloy systems [32, 33]. One explanation theorized from ab initio calculations on $\Sigma 5(310)$ GB is that boron improves grain boundary cohesion by a strong hybridization between the Fe $s,d$-states and the B $s,p$-states. In contrast, the hydrogen atom exhibits an ion-like character in metals [34, 35].
@@
While the localization and the quantification of hydrogen in metals and alloys remain experimentally challenging, different methods have been developed to detect this element in materials [36]. One of them is thermal desorption spectroscopy (TDS), which allows to quantify the amount of hydrogen located in the microstructure and estimates the trapping energy between hydrogen and different types of crystalline defects [37, 38, 39, 40, 41, 42, 4]. These defects acting as traps for hydrogen are usually identified from the TDS spectra either with complementary experiments [4], simulations [43], or by optimizing the microstructure to favor the formation of certain types of traps [38, 44]. However, discrepancies in binding/interaction energies between hydrogen and crystalline defects have been reported in the literature for martensitic steels [39] or even $\alpha$-iron [45] depending on the method used. For instance, the interaction energy of hydrogen in a single vacancy has been estimated from diffusion analysis around -0.55 eV [46] while ab initio calculations estimated this energy between -0.30 eV and -0.25 eV [47]). Such energy difference makes it difficult to conclude if this crystalline defect in $\alpha$-iron acts as a reversible or an irreversible trap for hydrogen, which is also observed for grain boundaries in martensitic steels [39].
@@
The present work focuses on the segregation of hydrogen in boron-doped steel and boron-free steel, respectively. After a careful investigation of the microstructure of both steels, they have been analyzed through TDS measurements. Then, ab initio calculations have been performed to understand the interaction between boron and hydrogen at grain boundaries. Finally, the equilibrium concentration of hydrogen has been evaluated from TDS measurements using the Langmuir-McLean approximation to determine the hydrogen partitioning in the microstructure when the two steels are hydrogen-charged.
@@ 2.2 Microstructure evolution with boron addition
Figure 1: IPF maps of a) LC and c) B-LC, and d) its corresponding reconstructed parent grains. The SGBs having a dislocation-like microstructure (with a misorientation lower than 8°) are represented in red and the MMBs with the KS variant misorientation and PAGB are represented in black in figs. a and c. The PAGBs are also represented in black in figs. b and d.
Further, three categories of grain boundaries have been sorted for LC and B-LC: (i) prior austenite grain boundaries that do not follow KS OR, (ii) martensite-martensite boundaries with a relevant KS OR, and (iii) subgrain boundaries (SGBs) having a misorientation lower than 8°. Table 1 presents the length of these three types of boundaries ($l_{PAGB}$, $l_{MMB}$, and $l_{SGB}$) for both steels. The similar percentage of all lengths suggests a similar microstructure with boron incorporation and most of the grain boundaries are SGBs and MMBs (more than 90% for LC and B-LC).
Table 1: Summary of the length distribution of PAGBs ($l_{PAGB}$), martensite-martensite boundaries following KS OR ($l_{MMB}$), and subgrain boundaries ($l_{LAGB}$) in LC and B-LC.
@@
The synchrotron X-ray measurements show similar two-dimensional diffractograms for LC and B-LC (figs. 2.a and 2.b, respectively). Peaks from the martensite microstructure and the retained austenite are obtained during the measurements when the circular integration of the two-dimensional diffractogram is presented in logarithmic scale. This austenite phase was not detected during the EBSD analyses, most likely because it is unstable and transforms to martensite when it reaches the free surfaces during metallographic sample preparation. When the synchrotron X-ray diffraction profiles are integrated, the contribution related to the austenite phase becomes clearer (fig. 2.c).
@@
Fig 2.d presents the volume fraction of the $\gamma$ phase ($f_{\gamma}$), which is 2.26$\pm$0.04 % for both steels and is deduced by integrating peaks related to the martensite and austenite phases. This method was chosen over the Rietveld refinement analysis approach because the latter neglects fluctuations that are observed at 4.5° and 7.5° (highlighted with black arrows in fig. 2.c). These fluctuations are related to carbides, which could be either cementite, $\eta$-carbides, or $\varepsilon$-carbides according to the literature [56] that some are observed through ECCI experiments (fig 10.d). The average lattice strain from the austenite phase has been calculated using [57]:
$\varepsilon^{ave}_{\gamma}=\frac{a_{\gamma}-a_{\gamma,0}}{a_{\gamma,0}},$ (1)
with $a_{\gamma}$ and $a_{\gamma,0}$, being the lattice constant measured from synchrotron X-ray diffraction and the lattice constant in the unstrained state, respectively. The unstrained lattice constant has been estimated from the carbon content in the material using the following empirical equation: $a_{\gamma,0}=0.35965+0.0028C_{C}$[58] and is 0.36007 nm. From eq. (1) a large lattice strain of -0.275% is deduced. The equivalent hydrostatic stress $\sigma_{hydro}^{\gamma}$ can subsequently be calculated using [59]:
$\sigma_{hydro}^{\gamma}=\frac{E}{1-2\nu}\varepsilon^{ave}_{\gamma},$ (2)
with $E$ and $\nu$ the Young’s modulus and Poisson’s ratio, respectively. When the Young’s modulus and Poisson’s ratio are chosen at 172 GPa, and 0.3 [59], a compressive hydrostatic stress $\sigma_{hydro}^{\gamma}$ of 1.2 GPa is deduced.
@@
Finally, correlative TKD-APT experiments have been performed to characterize the boron segregation at PAGBs. A site-specific lift-out has been performed for localizing PAGB from LC and B-LC microstructures, following the procedure presented in previous work [29] before the TKD observations. Figs 3.a and 3.d provide the TKD maps of LC and B-LC before the APT measurements. The orientation relationship of both GBs has first been checked to confirm that they are PAGBs. For LC, the misorientation angle/axis between the grains is 59°/[0.542 -0.497 -0.678] and is 44°/[-0.743 0.656 -0.131] for B-LC, which are the other grain is parallel to a $[9,4,10]$ direction. These orientation relationships do not follow the KS OR [62, 54], thus they can be considered as a PAGB.
@@
Figure 3: Correlative TKD-APT analysis of PAGBs in LC (a-c) and B-LC (d-f) steels. a) IPF map of LC with the three-dimensional reconstruction of the APT needle showing b) boron (red atoms) and c) carbon (blue atoms). d) IPF map of B-LC with the three-dimensional reconstruction showing e) boron (orange atoms) and f) carbon (green atoms). g) Corresponding one-dimensional composition profile in PAGBs.
@@
Figs 3.b, c present three-dimensional reconstructions showing only boron and carbon atoms for LC while figs. 3.e, f present three-dimensional reconstructions of the same elements in B-LC. Single events have been separated from multiple events to improve the visualization of boron and carbon in the reconstructed data. This data filtering has been shown to provide an improved signal-to-background ratio for species in low concentration in Fe-based materials [63]. More details are given at B, showing an improved quantification of boron in an interface obtained from martensitic transformations. A pronounced segregation of both carbon and boron atoms is observed into the GB for LC and B-LC, in agreement with the literature [64, 29, 26]. Figs 3.g plot the resulting composition profile of the interface marked by a blue cylinder in the corresponding APT reconstruction shown inset. While the concentration of carbon is similar for both interfaces (maximum concentration around 2.5 at.%), the maximum boron concentration is 0.12 at.for LC and higher than 1.2 at.for B-LC. This difference highlighted the pronounced segregation of boron in PAGBs for B-LC, which is not observed for LC, which contains almost no boron.
@@
In this section, we described the microstructure obtained in both LC and B-LC after quenching from 1100°C in He gas (cooling rate of $\sim$200 K.s-1). A similar microstructure is observed for LC and B-LC with an average grain size around 13 $\mu$m (when using the grain size of spherical grains) for both steels with the appearance of carbides, confirmed through synchrotron X-ray experiments. Additionally, these experiments show that 2 vol.of retained austenite is observed in these steels. This fraction of retained austenite is lower than the fraction deduced in medium carbon steel reported in the literature (which is 5.5%) [65] due to the difference in carbon and silicon content. This phase should be located at lath martensite boundaries according to the literature [65, 66]. The dislocation density in the martensite phase has been estimated from the synchrotron X-ray experiments as 1$\times$1015 m-2, and agrees with measurements performed in the literature [62]. In the austenite phase, a higher dislocation density has been estimated and is 3$\times$ 1015 m-2. Finally, the APT measurements correlated with TKD analysis reveal a major change in behavior between boron-doped and boron-free steels, which is the segregation of boron to the PAGBs. At these interfaces, the maximal boron concentration is $\sim$1 at.but can increase up to 8 at.according to previous work [29]. In addition, previous secondary ion mass spectroscopy analyses had demonstrated that the segregation at other martensite boundaries is minor [29]. Since it is expected to observe the same solute concentration for LC and B-LC at these interfaces, the trapping behavior towards hydrogen should be similar at these interfaces. Further, hydrogen is inserted into these microstructures and detected using TDS measurements. According to the microstructure characterization, hydrogen could diffuse and be potentially trapped in vacancies, dislocations (mostly screw dislocations, in their cores and their distortion fields), grain boundaries (martensitic laths, block and packet boundaries, PAGBs), precipitates (carbides), and in the austenite phase.

### v3

@@ Abstract
The interaction between boron and hydrogen at grain boundaries has been investigated experimentally and numerically in boron-doped and boron-free martensitic steels using thermal desorption spectrometry (TDS) and ab initio calculations. The calculations show that boron and hydrogen are attracted to grain boundaries but boron can repel hydrogen. This behavior has also been observed using TDS measurements, with the disappearance of one peak when boron is incorporated into the microstructure. Additionally, the microstructure of both steels has been studied through electron backscattered diffraction, synchrotron X-ray measurements, and correlative transmission Kikuchi diffraction-atom probe tomography measurements. While they have a similar grain size, grain boundary distribution, and dislocation densities, pronounced boron segregation into PAGBs is observed for boron-doped steels. It indicates that boron in PAGBs is responsible for the disappearance of the TDS peaks for the boron-doped steel. Then, the equilibrium hydrogen concentration in different trapping sites has been evaluated using the Langmuir-McLean approximation. This thermodynamic model shows that the distribution of hydrogen is identical for all traps when the total hydrogen concentration is low for boron-free steel. However, when it increases, traps of the lowest segregation energies (mostly PAGBs) are firstly saturated, which promotes failure initiation at this defect type. This finding partially explains why PAGBs are the weakest microstructure feature when martensitic steels are exposed to hydrogen-containing environments.
@@ 1 Introduction
Hydrogen is a promising carrier molecule for a low-carbon economy and many corresponding industrial projects are under development to use it as an energy carrier or to develop sustainable processes like the direct reduction of iron ore [1, 2, 3, 4, 5]. It is also known that hydrogen easily diffuses in metals and alloys inducing premature failure of engineering structures, a phenomenon called as hydrogen embrittlement (HE). HE is well-known and has been described by numerous models in the literature and thoroughly reviewed in refs. [6, 7, 8, 9]. However, these models provide only a phenomenological description of HE (e.g.: hydrogen enhancing the formation of vacancies [10, 11, 12], localizing plasticity [13, 14, 15], reducing the cohesive interface energy of the system [16, 17] and so on) because the direct observation of hydrogen at the atomic scale is quite challenging given the difficulty to locate hydrogen in alloys experimentally [18, 19, 20].
@@
Steel is one of the candidate materials envisaged to contain and transport hydrogen due to its high mechanical strength, low price, flexibility to form different microstructures from its chemical composition, and the variety of processes that can be applied to the alloys [21]. However, it is also subject to HE [22]. Different processes have been developed to reduce HE of steels [23, 24], including reducing the grain size [25], which is efficient up to a limit, and also modifies other properties of the alloy, potentially to a point of being unsuitable for the envisaged application. Steel composition features have also been modified, for instance, to form precipitates and reduce the negative impact of hydrogen by trapping the solute [26, 27]. However, once traps inherent to the precipitates, including interfaces with the matrix or sites inside the precipitates, are saturated with hydrogen, they become inefficient for further HE mitigation specifically in typical engineering loading scenarios, which are often characterized by permanent hydrogen exposure. More recently, the implementation of chemically heterogeneous Mn profiles in the austenite phase of medium Mn-steels has also been proposed, improving the ductility of the alloys in hydrogen [28]. However, such treatment is only possible for a relatively narrow range of Mn-steel compositions.
@@
In martensitic steel, hydrogen reduces the cohesion of prior austenite grain boundaries (PAGBs), which is the weakest microstructure feature in steel in a hydrogen environment [29, 30]. A recently proposed solution to mitigate HE in martensitic steels is to favor carbon segregation [30]. In the present work, boron, which is another interstitial solute segregating in PAGBs, should improve the cohesion of the crystalline defect and could mitigate hydrogen embrittlement. It is an alloying element often incorporated in steel because it improves the hardenability and delays the formation of ferrite [31]. However, the segregation mechanism of boron remained elusive until recently, when it was described using equilibrium segregation and considering the precipitation of borides or boro-carbides at GBs [32, 33]. It has also been observed recently that boron can improve the resistance to HE of steels [34], which is also observed in other alloy systems [35, 36]. One explanation theorized from ab initio calculations on $\Sigma 5(310)$ GB is that boron improves grain boundary cohesion by a strong hybridization between the Fe $s,d$-states, and the B $s,p$-states. In contrast, the hydrogen atom exhibits an ion-like character in metals [37, 38] but experimental evidence of this behavior is still missing in the literature.
@@
While the localization and the quantification of hydrogen in metals and alloys remain experimentally challenging, different methods have been developed to detect this element in materials [39]. One of them is thermal desorption spectroscopy (TDS), which allows to quantify the amount of hydrogen located in the microstructure and estimates the trapping energy between hydrogen and different types of crystalline defects [40, 41, 42, 43, 44, 45, 5]. These defects acting as traps for hydrogen are usually identified from the TDS spectra either with complementary experiments [5], simulations [46], or by optimizing the microstructure to favor the formation of certain types of traps [41, 47]. However, discrepancies in binding/interaction energies between hydrogen and crystalline defects have been reported in the literature for martensitic steels [42] or even $\alpha$-iron [48] depending on the method used. For instance, the interaction energy of hydrogen in a single vacancy has been estimated from diffusion analysis around -0.55 eV [49] while ab initio calculations estimated this energy between -0.30 eV and -0.25 eV [50]). Such energy difference makes it difficult to conclude if this crystalline defect in $\alpha$-iron acts as a reversible or an irreversible trap for hydrogen, which is also observed for grain boundaries in martensitic steels [42]. Additionally, martensitic steels contain a complex microstructure with different types of grain boundaries from austenitisation and martensite transformation. When TDS measurements are performed on hydrogen-charged ultra-low-carbon and high-carbon steels, several temperature ranges have been identified for the desorption of hydrogen from grain boundaries (average and high-angle grain boundaries [51, 47, 52]), making difficult to interpret the hydrogen desorption rate of martensite steels.
@@
The present work aims to clarify these points by focusing the study on the segregation of hydrogen in boron-doped and boron-free steels. After a careful investigation of the microstructure of both steels, they have been analyzed through TDS measurements. Then, ab initio calculations have been performed to understand the interaction between boron and hydrogen at grain boundaries. Finally, the equilibrium concentration of hydrogen has been evaluated from TDS measurements using the Langmuir-McLean approximation to determine the hydrogen partitioning in the microstructure when the two steels are hydrogen-charged.
@@ 2.2 Microstructure evolution with boron addition
Figure 1: In plane IPF maps of a) LC and c) B-LC, and d) its corresponding reconstructed parent grains. The SGBs having a dislocation-like microstructure (with a misorientation lower than 8°) are represented in red and the MMBs with the KS variant misorientation and PAGB are represented in black in figs. a and c. The PAGBs are also represented in black in figs. b and d.
Further, three categories of grain boundaries have been sorted for LC and B-LC: (i) prior austenite grain boundaries that do not follow KS OR, (ii) martensite-martensite boundaries (MMBs) with a relevant KS OR, and (iii) subgrain boundaries (SGBs) having a misorientation lower than 8°. Table 1 presents the length distribution in the fixe of these three types of boundaries ($l_{PAGB}$, $l_{MMB}$, and $l_{SGB}$) for both steels. Both microstructures contain mostly grain boundaries from martensite transformation (more than 90% of GBs are SGBs and MMBs for LC and B-LC). The distribution of grain boundary types in LC and B-LC indicates a similar microstructure with boron incorporation and suggests that boron has a minor impact on the grain boundary distribution.
Table 1: Summary of the length distribution of PAGBs ($l_{PAGB}$), MMBs following KS OR ($l_{MMB}$), and SGBs ($l_{LSGB}$) in LC and B-LC. They are the sum of segments’ length from the map shown in fig. 1.
@@
The synchrotron X-ray measurements show similar two-dimensional diffractograms for LC and B-LC (figs. 2.a and 2.b, respectively). Peaks from the martensite microstructure and the retained austenite are obtained during the measurements when the circular integration of the two-dimensional diffractogram is presented in logarithmic scale. This austenite phase should be located at lath martensite boundaries according to the literature for medium carbon steels [61, 62] and was not detected during the EBSD analyses. It is either because it is unstable and transforms to martensite when it reaches the free surfaces during metallographic sample preparation or the thickness of the austenite layer is too thin to be detected by EBSD. When the synchrotron X-ray diffraction profiles are integrated, the contribution related to the austenite phase becomes clearer (fig. 2.c).
@@
Fig 2.d presents the volume fraction of the $\gamma$ phase ($f_{\gamma}$), which is 2.26$\pm$0.04 % for both steels and is deduced by integrating peaks related to the martensite and austenite phases. This method was chosen over the Rietveld refinement analysis approach because the latter neglects fluctuations that are observed at 4.5° and 7.5° (highlighted with black arrows in fig. 2.c). These fluctuations are related to carbides, which could be either cementite, $\eta$-carbides, or $\varepsilon$-carbides according to the literature [63] that some are observed through ECCI experiments (fig 10.d).
@@
Finally, correlative TKD-APT experiments have been performed to characterize the boron segregation at PAGBs and MMBs. A site-specific lift-out has been performed for localizing the different grain boundaries from LC and B-LC microstructures, following the procedure presented in previous work [32] before the TKD observations. Figs 3.a and 3.d provide the TKD maps of PAGBs and figs 3.h and 3.k of MMBs from LC and B-LC before the APT measurements, respectively. The orientation relationship of the different GBs has first been checked to confirm the nature of each boundary. For PAGB, the misorientation angle/axis between the grains is 44°/[0.526 -0.316 -0.789] for the TKD map of fig 3.a and is 59°/[0.539 -0.539 0.647] for the TKD map of fig 3.h. The latter orientation is close to the variant V1 obtained from the KS OR (60°/[0.577 -0.577 0.577]) [66, 59], indicating that it is a MMB while the former is a PAGBs. For B-LC, it is 44°/[-0.743 0.656 -0.131] for the TKD map of fig 3.d and it is 52°/[-0.734 -0.587 0.342] for the TKD map of fig 3.k. The latter grain boundary of B-LC is close to the variant V17 obtained from the KS OR (51.7°/[-0.659 0.363 -0.59]) [66, 59], it suggests that it is a MMB and while the grain boundary of fig 3.d can be considered as a PAGB.
@@
Figure 3: Correlative TKD-APT analysis of PAGBs (a-g) and martensite-martensite boundaries (h-n) in LC and B-LC steels. a) (and h) IPF map of LC with the three-dimensional reconstruction of the APT needle showing b) (and i) boron (red atoms) and c) (and j) carbon (blue atoms). d) (and k) IPF map of B-LC with the three-dimensional reconstruction showing e) (and l) boron (orange atoms) and f) (and m) carbon (green atoms). g) (and n) Corresponding one-dimensional composition profile in PAGBs (in MMBs).
@@
Figs 3.b, c present three-dimensional reconstructions showing only boron and carbon atoms for LC while figs. 3.e, f present three-dimensional reconstructions of the same elements in B-LC. Single events have been separated from multiple events to improve the visualization of boron and carbon in the reconstructed data. This data filtering has been shown to provide an improved signal-to-background ratio for species in low concentration in Fe-based materials [67]. More details are given at B, showing an improved boron quantification in an interface obtained from martensitic transformations. Figs 3.i, j and figs. 3.l, m present a similar reconstruction but for a MMB for LC and B-LC, respectively. For all reconstructions, the grain boundary can be localized from the pronounced carbon segregation. Figs 3.g and 3.n plot the resulting composition profile of the interface marked by a blue cylinder in the corresponding APT reconstruction shown inset. While the concentration of carbon is similar for all interfaces (maximum concentration around 2.5 at.%), the maximum boron concentration is less than 0.1 at.for LC and higher than 1.2 at.for B-LC in PAGBs. However, this pronounced segregation is not observed in MMBs of B-LC. This difference highlighted the pronounced segregation of boron in PAGBs for B-LC, which is not observed for LC, or in MMBs of B-LC, which contains almost no boron.
@@
In this section, we described the microstructure obtained in both LC and B-LC after quenching from 1100°C in He gas (cooling rate of $\sim$200 K.s-1). A similar microstructure is observed for LC and B-LC with an average grain size around 13 $\mu$m (when using the grain size of spherical grains) for both steels with the appearance of carbides, confirmed through synchrotron X-ray experiments. Additionally, these experiments show that 2 vol.of retained austenite is observed in these steels. This fraction of retained austenite is lower than the fraction deduced in medium carbon steel reported in the literature (which is 5.5%) [61]. The dislocation density in the martensite phase has been estimated from the synchrotron X-ray experiments as 1$\times$1015 m-2, and agrees with measurements performed in the literature [66]. In the austenite phase, a higher dislocation density has been estimated and is 3$\times$ 1015 m-2. Finally, the APT measurements correlated with TKD analysis reveal a major change in behavior between boron-doped and boron-free steels, which is the segregation of boron to the PAGBs. At these interfaces, the maximal boron concentration is $\sim$1 at.but can increase up to 8 at.according to previous work [32]. In addition, previous secondary ion mass spectroscopy analyses had demonstrated that the segregation at other martensite boundaries is minor [32]. Since it is expected to observe the same solute concentration for LC and B-LC at these interfaces, the trapping behavior towards hydrogen should be similar at these interfaces. Further, hydrogen is inserted into these microstructures and detected using TDS measurements. According to the microstructure characterization, hydrogen could diffuse and be potentially trapped in vacancies, dislocations from martensite transformation, grain boundaries (martensitic laths, block and packet boundaries, PAGBs), precipitates (carbides), and in the austenite phase.